What is the magnitude of the force a 1.5 x 10-6C charge exerts on a 3.2 x 10 -4C charge located 2.5 m away?


1
Expert's answer
2021-11-21T17:16:45-0500

We know that

Coulamb force

"F=\frac{kq_1q_2}{r^2}"

"F=\frac{9\times10^9\times1.5\times\times10^{-6}\times3.2\times10^{-4}}{2.5^2}"

"F=0.6912N"